OpenStudy (anonymous):

Help, Will medal and fan! George and carman went on a bicycle ride. they took a bus to their starting point, and they biked the rest. they traveled 300 kilometers in total, and the miked 70 kilometers more than they were bussed. How many kilometers did they bike?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@natsent06 @campbell_st

OpenStudy (princeharryyy):

let they bused x kms then biked = x+70 => x+70+x = 300 x=115km xo bike = 115 + 70 =185 km

OpenStudy (princeharryyy):

@ALG3BRUHH69

OpenStudy (anonymous):

thats what i got too, but it says thats incorrect

OpenStudy (princeharryyy):

den i cant do nything in that